# Deep Image Composition Meets Image Forgery

### v1

@@ 3 Proposed Framework
Figure 4: Comparison of proposed framework
@@ 4 Evaluation
In order to measure the quality of the dataset we produced, the Early Fusion image manipulation detection model was used to make predictions. The SOTA detection model achieved 71.899% classification success on 24964 images in our dataset. Compared to other datasets "see Table 2", it achieved a lower success rate, which shows how high quality and difficult the images we produced were to detect. In addition, to test our hypothesis that image composition techniques help hide manipulation, we also had the model predict original OPA images. We found that the success rate was 78.38% when image composition techniques were not applied to the same image composition. Unlike double compression at low quality level, adding noise and blurring the image, our proposed framework reduces the manipulation detection rate by 6.48% without affecting the image quality.
Columbia Casiav1+ DSO-1 OPA Ours
Classification Accuracy(%) 96.2 84.5 93.5 78.38 71.899
@@ 5 Future Work
11:end for

### v2

@@ 3 Proposed Framework
Figure 4: Comparison of proposed framework(zoom foreground objects’ edges for details)
@@ 4 Evaluation
In order to measure the quality of the dataset we produced, the Early Fusion image manipulation detection model was used to make predictions. Images were resized to 512x512 for convenience before making predictions. We taught it would not result in any accuracy drop because a very similar image manipulation detection study Trufor is not efected by image resizing [21]. With resizing the SOTA detection model achieved 71.899% classification success on 24964 images in our dataset. In addition, to test our hypothesis that image composition techniques help hide manipulation, we also had the model predict original OPA images. We found that the success rate was 78.38% when image composition techniques were not applied to the same image composition.
We did also check the accuracy of the SOTA model on other datasets. Despite the fact that image resizing (without changing image aspect ratio) data augmentation was employed during the training of the SOTA model, with resizing images to 512x512 resolution a drastic classification accuracy drop had been observed. One possible reason to this accuracy drop may be because of resizing by changing the image aspect ratio because the SOTA model has data augmentation without changing aspect ratio.
After this discovery, we did tested our dataset and corresponding OPA images without resizing. The detection model had 86.84% accuracy on our dataset and 87.74% on OPA images, in other words no significant accuracy change observed. Our guess on the reason for resizing without keeping aspect ratio affect our hypothesis test is the removal of compression and noise artifacts. Consequently our dataset, that has less spatial domain artifacts and color inconsistency, is harder to detect without noise and compression artifacts.
Compared to other datasets "see Table 2", it achieved a lower success rate, which shows how high quality and difficult the images we produced were to detect. Unlike double compression at low quality level, adding noise and blurring the image, our proposed framework reduces the manipulation detection rate by 6.48% without affecting the image quality.
In splicing images, as a rule, the ratio of the area of the foreground object to the area of the composite image should not exceed 50%. If it exceeds this, there will be confusion when identifying foreground and background regions. In the dataset we created, there are very few images where this ratio exceeds 50% "see Figure 5". Since this is a POC study, it can be neglected.
Figure 5: Histogram of manipulated area ratios
Columbia Casiav1+ DSO-1 OPA Ours OPA(512x512) Ours(512x512)
Classification Accuracy(%) 96.2 84.5 93.5 87.74 86.94 78.38 71.899
@@ 5 Future Work
11:     Employ different kind of post processing attacks
12:end for